OpenStudy (anonymous):

Simplify by removing the factors of 1. S^2-81/S^2-18s+81 please show the steps.

OpenStudy (anonymous):

\[\frac{s^2-81}{s^2-18s+81}=\frac{(s-9)(s+9)}{(s-9)(s-9)}\]

OpenStudy (anonymous):

so when you "remove a the factors of 1" aka cancel, you get \[\frac{s+9}{s-9}\]

OpenStudy (anonymous):

I know this is dumb but what is the 1 factor

OpenStudy (anonymous):

what they mean is this

OpenStudy (anonymous):

\[\frac{(s-9)(s+9)}{(s-9)(s-9)}=\frac{s-9}{s-9}\times \frac{s+9}{s-9}=1\times \frac{s+9}{s-9}\]

OpenStudy (anonymous):

so this odd wording just means that \[\frac{s-9}{s-9}=1\] so you can remove it

OpenStudy (anonymous):

it is not a dumb question, it is just worded funny. who wrote this?

OpenStudy (anonymous):

my professor

OpenStudy (anonymous):

figures

OpenStudy (anonymous):

lol yea It has been one confusing class I can tell you this much.

OpenStudy (anonymous):

good luck! hope this helped

OpenStudy (anonymous):

Yes thanks I am sure i will be posting more hahaha.